Allen bought 20 stamps at the post office in 37¢ and 20¢ denominations. If the total cost of the stamps was $7.06, how many 37¢ stamps did Allen buy?
Answer by stanbon(75887)

Allen bought 20 stamps at the post office in 37¢ and 20¢ denominations. If the total cost of the stamps was $7.06, how many 37¢ stamps did Allen buy?
---------------
Let # of 20 cent stamps be "x" ; Value of these is 20x cents
---------
# of 37 cent stamps is "20-x" ; Value of these is 37(20-x)=740-37x cents
-------------------
EQUATION:
value + value = 706 cents
20x + 740-37x = 706
-17x=-34
x=2 (# of 20 cent stamps)
20-x = 18 (# of 37 cent stamps)
---------
Cheers,
Stan H.